# Positive definiteness and Fell bundles over discrete groups

### v1

@@ 3.2 Completely positive maps
where $w_{i,j}:=M(\lambda^{\mathcal{A}}_{g_{i}}(a_{i})^{*}\lambda^{\mathcal{A}}_{g_{j}}(a_{j}))\in C_{r}^{*}({\mathcal{B}})$.
@@
Using [27, Proposition 17.12] and setting $\Omega:=\big{(}j^{\mathcal{B}}_{\varphi(g_{1})}(b_{1})^{*},\ldots,j^{\mathcal{B}}_{\varphi(g_{n})}(b_{n})^{*}\big{)}\in(\ell^{2}({\mathcal{B}}))^{n}$, we therefore get that
@@
Thus, an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is reduced when there exists a (uniquely determined) linear bounded map $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ satisfying that $M_{T}\circ\iota^{\mathcal{A}}=\iota^{\mathcal{B}}\circ\phi_{T}$, that is,
$M_{T}\Big{(}\sum_{g\in G}\lambda^{\mathcal{A}}_{g}(f(g))\Big{)}=\sum_{g\in G}\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(f(g)))\quad\text{ for all }~f\in C_{c}({\mathcal{A}}),$
[…]
Similarly, an ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T=(T_{g})_{g\in G}$ is full when there exists a (uniquely determined) linear bounded map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ satisfying that $\Phi_{T}\circ\kappa^{\mathcal{A}}=\kappa^{\mathcal{B}}\circ\phi_{T}$, that is,
$\Phi_{T}\Big{(}\sum_{g\in G}\,\widehat{j}^{\mathcal{A}}_{g}(f(g))\Big{)}=\sum_{g\in G}\,\widehat{j}^{\mathcal{B}}_{\varphi(g)}(T_{g}(f(g)))\quad\text{ for all }~f\in C_{c}({\mathcal{A}}),$
[…]
We note that if $T=(T_{g})_{g\in G}$ is a reduced ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map such that $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ is completely positive, then $T$ is positive definite. Indeed, this follows from Proposition 3.7 (with $M=M_{T}$), since
$T_{g}(a)=E^{\mathcal{B}}_{\varphi(g)}\big{(}\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\big{)}=E^{\mathcal{B}}_{\varphi(g)}\big{(}M_{T}(\lambda^{\mathcal{A}}_{g}(a))\big{)}$
for every $g\in G$ and $a\in A_{g}$.
We also note that if $T$ is a full ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map such that $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ is completely positive, then $T$ is positive definite. Indeed, for every $g\in G$ and $a\in A_{g}$, we have
$\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))=\Lambda^{\mathcal{B}}(\,\widehat{j}^{\,{\mathcal{B}}}_{\varphi(g)}(T_{g}(a)))=\Lambda^{\mathcal{B}}(\Phi_{T}(\,\widehat{j}^{\,{\mathcal{A}}}_{g}(a))),$
thus
$T_{g}(a)=E^{\mathcal{B}}_{\varphi(g)}\big{(}\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\big{)}=E^{\mathcal{B}}_{\varphi(g)}\big{(}\Lambda^{\mathcal{B}}(\Phi_{T}(\widehat{j}^{\,{\mathcal{A}}}_{g}(a)))\big{)},$
and the assertion follows from Proposition 3.8 (with $\Phi=\Phi_{T}$).
We will now discuss whether the converse of these two assertions hold. We will first show that it does in the full case. To obtain a reduced version, we will have to assume that the kernel of $\varphi$ is amenable. Such a condition is known to be necessary and sufficient when considering the extendibility of the homomorphism $\varphi:G\to H$ to a *-homomorphism from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ (cf. [11, Corollary 8.C.15]). Note here that a completely positive map from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ which extends $\varphi$ has to be a *-homomorphism, so one gets from the result cited above that $\varphi$ extends to a completely positive linear map from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ if and only if ${\rm ker}(\varphi)$ is amenable.
@@
By [18, Corollary A.8], there exists a completely positive map $\tilde{\phi}:C^{*}({\mathcal{A}})\to B({\mathcal{H}})$ such that
@@
Note that if ${\mathcal{A}}$ and ${\mathcal{B}}$ are amenable in the sense of Exel [26, 27], i.e., their full and reduced cross-sectional $C^{*}$-algebras are canonically *-isomorphic, respectively, then we get as a corollary to Theorem 3.10 that $T$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map if and only if $T$ is reduced and $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ is completely positive. When ${\mathcal{A}}$ or ${\mathcal{B}}$ are not amenable, this equivalence is also true under the assumption that ${\rm ker}(\varphi)$ is an amenable subgroup of $G$. In order to show this, we will first generalize [18, Proposition 4.8].
@@
Then there is a completely positive linear map $\psi^{\prime}:C^{*}_{r}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ such that for all $g\in G$
$\psi^{\prime}\big{(}\lambda^{\mathcal{A}}_{g}(A_{g})\big{)}\subseteq\,\lambda^{\mathcal{B}}_{\varphi(g)}(B_{\varphi(g)})$ (3)
and
@@
By Stinespring’s dilation theorem for completely positive maps (see e.g. [46, Theorem 5.6]), there exist a Hilbert space ${\mathcal{K}}$, a *-homomorphism $\widehat{\psi}:C^{*}({\mathcal{A}})\to B({\mathcal{K}})$ and a bounded operator $V:{\mathcal{H}}\to{\mathcal{K}}$ such that $\psi(x)=V^{*}\,\widehat{\psi}(x)\,V$ for all $x\in C^{*}({\mathcal{A}})$. Let $\widehat{\pi}=(\widehat{\pi}_{g})_{g\in G}$ denote the representation of ${\mathcal{A}}$ on ${\mathcal{K}}$ associated to $\widehat{\psi}$, so that
@@
By Fell’s absorption principle for Fell bundles [27, Proposition 18.4], the integrated forms of the representations $\widehat{\pi}\otimes\lambda^{G}=(\widehat{\pi}_{g}\otimes\lambda^{G}_{g})_{g\in G}$ and $\pi\otimes\lambda^{H}=(\pi_{h}\otimes\lambda^{H}_{h})_{h\in H}$ factor through $C^{*}_{r}({\mathcal{A}})$ and $C^{*}_{r}({\mathcal{B}})$, respectively. Let $\psi_{\mathcal{A}}:C^{*}_{r}({\mathcal{A}})\to B({\mathcal{K}}\otimes\ell^{2}(G))$ and $\psi_{\mathcal{B}}:C^{*}_{r}({\mathcal{B}})\to B({\mathcal{H}}\otimes\ell^{2}(H))$ denote the induced *-homomorphisms. In fact, using [27, Corollary 18.5], we may regard $\Psi_{\mathcal{A}}$ as a map from $C^{*}_{r}({\mathcal{A}})$ into $B({\mathcal{K}})\otimes_{\rm min}C_{r}^{*}(G)$ and $\Psi_{\mathcal{B}}$ as a map from $C^{*}_{r}({\mathcal{B}})$ into $B({\mathcal{H}})\otimes_{\rm min}C_{r}^{*}(H)$, satisfying that
@@
for all $g\in G$, $a\in A_{g}$, $h\in H$ and $b\in B_{h}$. Note that $\psi_{\mathcal{B}}$ is faithful since $\pi_{e}$ is assumed to be so, cf. [27, Proposition 18.4].
Now, as shown in [11, Corollary 8.C.15], $\varphi$ extends to a *-homomorphism $\varphi_{*,r}:C_{r}^{*}(G)\to C_{r}^{*}(H)$ satisfying that $\varphi_{*,r}(\lambda_{g}^{G})=\lambda^{H}_{\varphi(g)}$ for all $g\in G$.
@@
$~\Phi(\lambda^{\mathcal{A}}_{g}(a))$ $=(\widetilde{V}\otimes_{\rm min}\varphi_{*,r})(\psi_{\mathcal{A}}(\lambda^{\mathcal{A}}_{g}(a)))$
$=(\widetilde{V}\otimes_{\rm min}\varphi_{*,r})(\widehat{\pi}_{g}(a)\otimes\lambda^{G}_{g})$
$=V^{*}\widehat{\pi}_{g}(a)V\otimes\lambda^{H}_{\varphi(g)}$
$=V^{*}\widehat{\psi}(\widehat{j}^{\mathcal{A}}_{g}(a))V\otimes\lambda^{H}_{\varphi(g)}$
@@
Since $\psi_{\mathcal{B}}$ is faithful, we can define $\psi^{\prime}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ by $\psi^{\prime}=(\psi_{\mathcal{B}})^{-1}\circ\Phi$, which then satisfies $\psi^{\prime}(\lambda^{\mathcal{A}}_{g}(A_{g}))\subseteq\lambda^{\mathcal{B}}_{\varphi(g)}(B_{\varphi(g)})$, i.e., (3) holds. It is now straightforward to check that $\psi^{\prime}$ also satisfies (4). $\square$
@@
for $g\in G$ and $a\in A_{g}$. Since $\pi_{\varphi(g)}$ is injective on $B_{\varphi(g)}$, this implies that
$\phi^{\prime}(\lambda^{\mathcal{A}}_{g}(a))=\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\quad\text{for every }g\in G\text{ and }a~\in A_{g}.$
@@
Assume ${\rm ker}(\varphi)$ is not amenable and $T=(T_{g})_{g\in G}$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Then it may still happen that $T$ is reduced and $M_{T}$ is completely positive, cf. Example 4.1. Hence, in such a setting, it would be interesting to find some additional condition(s) ensuring that $T$ is reduced with $M_{T}$ completely positive.
@@
One could say that a completely positive linear map $\Psi:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ is $\varphi$-diagonal whenever
$\Psi(\,\widehat{j}^{\,{\mathcal{A}}}_{g}(A_{g}))\subseteq\,\,\widehat{j}^{\,{\mathcal{B}}}_{\varphi(g)}(B_{\varphi(g)})$
for all $g\in G$. Note that if $T$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map, then $\Phi_{T}$ is $\varphi$-diagonal. For a general completely positive linear map $\Phi:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$, letting $T^{\Phi}$ denote the positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map associated to it in Proposition 3.7, we get from Theorem 3.10 that the map $\Phi_{T^{\Phi}}$ is a $\varphi$-diagonal completely positive linear map (which could be called the $\varphi$-diagonalization of $\Phi$). Hence the map $\Phi\mapsto\Phi_{T^{\Phi}}$ is a projection from the space CP($C^{*}({\mathcal{A}}),C^{*}({\mathcal{B}})$) of completely positive linear maps onto the subspace of $\varphi$-diagonal maps. If the kernel of $\varphi$ is amenable, using Proposition 3.8 and Theorem 3.12, one obtains a similar projection map on $CP(C_{r}^{*}({\mathcal{A}}),C_{r}^{*}({\mathcal{B}}))$.
@@
Then any $S\in B({\mathcal{A}},\varphi,{\mathcal{B}})$ is a full ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map, which is also reduced if ${\rm ker}(\varphi)$ is assumed to be amenable. This follows readily from Theorems 3.10 and 3.12 after noticing that the map $T\mapsto\phi_{T}$ is linear. We thus get a linear map $T\mapsto\Phi_{T}$ from $B({\mathcal{A}},\varphi,{\mathcal{B}})$ into the space of completely bounded linear maps CB($C^{*}({\mathcal{A}}),C^{*}({\mathcal{B}})$), and a similar one into CB($C_{r}^{*}({\mathcal{A}}),C_{r}^{*}({\mathcal{B}})$) when ${\rm ker}(\varphi)$ is amenable. It is easy to see that the space $B({\mathcal{A}}):=B({\mathcal{A}},{\rm id}_{G},{\mathcal{A}})$ becomes a unital algebra w.r.t. natural composition, which could be called the Fourier-Stieltjes algebra of ${\mathcal{A}}$, in analogy with the case where ${\mathcal{A}}$ is the Fell bundle associated to a unital discrete twisted $C^{*}$-dynamical system, cf. [9].
@@
3.3 Categorical considerations
There is a category ${\tt Fell}$-${\tt PD}$ whose objects are Fell bundles over discrete groups, and an arrow from ${\mathcal{A}}=(A_{g})_{g\in G}$ to ${\mathcal{B}}=(B_{h})_{h\in H}$ is defined as a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map for some homomorphism $\varphi:G\to H$, with composition of arrows defined in the natural way. More formally, we define an arrow from ${\mathcal{A}}=(A_{g})_{g\in G}$ to ${\mathcal{B}}=(B_{h})_{h\in H}$ to be a pair $(\varphi,T)$, where $\varphi:G\to H$ is a homomorphism and $T$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map. Composition of an arrow $(\varphi,T)$ from ${\mathcal{A}}$ to ${\mathcal{B}}$ with an arrow $(\psi,S)$ from ${\mathcal{B}}$ to ${\mathcal{C}}=(C_{k})_{k\in K}$ is defined by
$(\psi,S)\circ(\varphi,T)=(\psi\varphi,R)$
where $R$ is the ${\mathcal{A}}$-$\psi\varphi$-${\mathcal{C}}$ bundle map given by $R:=(S_{\varphi(g)}T_{g})_{g\in G}$. Note here that $R$ is positive definite, hence that $(\psi\varphi,R)$ is an arrow in ${\tt Fell}$-${\tt PD}$. Indeed, as
$(\Phi_{S}\circ\Phi_{T})\Big{(}\widehat{j}^{\mathcal{A}}_{g}(b)\Big{)}=\Phi_{S}\Big{(}\widehat{j}^{\mathcal{B}}_{\varphi(g)}(T_{g}(b))\Big{)}=\widehat{j}^{\mathcal{C}}_{\psi(\varphi(g))}(S_{\varphi(g)}(T_{g}(b)))=\widehat{j}^{\mathcal{C}}_{(\psi\varphi)(g)}(R_{g}(b)))$
for all $g\in G\text{ and}~b\in B_{g}$, we get that $\Phi_{S}\circ\Phi_{T}$ is a completely positive linear map from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{C}})$ which extends $\phi_{R}$, and it follows from Theorem 3.10 that $R$ is positive definite.
For an object ${\mathcal{A}}$ in ${\tt Fell}$-${\tt PD}$, the identity arrow is defined as $({\rm id}_{G},{\rm id}_{\mathcal{A}})$ where ${\rm id}_{G}$ is the identity map on $G$ and ${\rm id}_{\mathcal{A}}=(({\rm id}_{\mathcal{A}})_{g})_{g\in G}$ with $({\rm id}_{\mathcal{A}})_{g}(a):=a$ for all $g\in G$ and $a\in A_{g}$. It is now straightforward to check that ${\tt Fell}$-${\tt PD}$ becomes a category.
The main point with introducing ${\tt Fell}$-${\tt PD}$ is that we get a functor $\mathcal{F}$ from the category ${\tt Fell}$-${\tt PD}$ to the category ${\tt C^{*}CP}$ whose objects are $C^{*}$-algebras and arrows are completely positive linear maps, by setting

### v2

@@ 3.2 Completely positive maps
where $w_{ij}:=M(\lambda^{\mathcal{A}}_{g_{i}}(a_{i})^{*}\lambda^{\mathcal{A}}_{g_{j}}(a_{j}))\in C_{r}^{*}({\mathcal{B}})$.
@@
Using [28, Proposition 17.12] and setting $\Omega:=\big{(}j^{\mathcal{B}}_{\varphi(g_{1})}(b_{1})^{*},\ldots,j^{\mathcal{B}}_{\varphi(g_{n})}(b_{n})^{*}\big{)}\in(\ell^{2}({\mathcal{B}}))^{n}$, we therefore get that
@@
In other words, $T$ is reduced precisely when there exists a (necessarily unique) linear bounded map $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ such that
$M_{T}\Big{(}\lambda^{\mathcal{A}}_{g}(a)\Big{)}=\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\,\,\text{ for all}~g\in G\text{ and}~a\in A_{g}.$
Similarly, $T$ is full precisely when there exists a (necessarily unique) linear bounded map $\Phi_{T}:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ such that
$\Phi_{T}\Big{(}\widehat{j}^{\mathcal{A}}_{g}(a)\Big{)}=\widehat{j}^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\,\,\text{ for all}~g\in G\text{ and}~a\in A_{g}.$
We note that if $T$ is reduced (resp. full) and $M_{T}$ (resp. $\Phi_{T}$) is completely positive, then $T$ is positive definite. Indeed, the first claim follows from Proposition 3.8 (with $M=M_{T}$), since
$T_{g}(a)=E^{\mathcal{B}}_{\varphi(g)}\big{(}\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\big{)}=E^{\mathcal{B}}_{\varphi(g)}\big{(}M_{T}(\lambda^{\mathcal{A}}_{g}(a))\big{)}\,\,\text{ for all}~g\in G\text{ and}~a\in A_{g},$
while the second one follows in a similar fashion from Proposition 3.9 (with $\Phi=\Phi_{T}$).
We will now discuss whether the converse of these last two assertions hold. We will first show that it does in the full case. To obtain a reduced version, we will have to assume that the kernel of $\varphi$ is amenable. Such a condition is known to be necessary and sufficient when considering the extendibility of the homomorphism $\varphi:G\to H$ to a *-homomorphism from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ (cf. [11, Corollary 8.C.15]). Note here that a completely positive map from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ which extends $\varphi$ has to be a *-homomorphism, so one gets from the result cited above that $\varphi$ extends to a completely positive linear map from $C_{r}^{*}(G)$ into $C_{r}^{*}(H)$ if and only if ${\rm ker}(\varphi)$ is amenable.
@@
By [18, Corollary A.8], there exists a completely positive map $\tilde{\phi}:C^{*}({\mathcal{A}})\to B({\mathcal{H}})$ (called the “integrated form” of $\phi$ in [18]) such that
@@
Note that if ${\mathcal{A}}$ and ${\mathcal{B}}$ are amenable in the sense of Exel [27, 28], i.e., their full and reduced cross-sectional $C^{*}$-algebras are canonically *-isomorphic, respectively, then we get as a corollary to Theorem 3.11 that $T$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map if and only if $T$ is reduced and $M_{T}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ is completely positive. When ${\mathcal{A}}$ or ${\mathcal{B}}$ are not amenable, this equivalence is also true under the assumption that ${\rm ker}(\varphi)$ is an amenable subgroup of $G$. In order to show this, we will first generalize [18, Proposition 4.8]. We will use the following terminology.
Definition 3.12.
A linear map $\Phi:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$ (resp. $\Psi:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$) will be called $\varphi$-diagonal whenever $\Phi(\,\widehat{j}^{\,{\mathcal{A}}}_{g}(A_{g}))\subseteq\,\,\widehat{j}^{\,{\mathcal{B}}}_{\varphi(g)}(B_{\varphi(g)})$ (resp. $\Psi\big{(}\lambda^{\mathcal{A}}_{g}(A_{g})\big{)}\subseteq\,\lambda^{\mathcal{B}}_{\varphi(g)}(B_{\varphi(g)})$) for all $g\in G$.
For example, if $T$ is a full (resp. reduced) ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map, then the associated map $\Phi_{T}$ (resp. $M_{T}$) is $\varphi$-diagonal. It is easy to see that $\{\Phi_{T}:T\text{ is a full ${\mathcal{A}}$-$\varphi$-${\mathcal{B}}$ bundle map}\}$ coincides with the set of $\varphi$-diagonal bounded linear maps from $C^{*}({\mathcal{A}})$ into $C^{*}({\mathcal{B}})$. A similar statement holds for $\varphi$-diagonal bounded linear maps from $C_{r}^{*}({\mathcal{A}})$ into $C_{r}^{*}({\mathcal{B}})$.
@@
Then there is a $\varphi$-diagonal completely positive linear map $\psi^{\prime}:C^{*}_{r}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ such that for all $g\in G$
@@
By Stinespring’s dilation theorem for completely positive maps (see e.g. [47, Theorem 5.6]), there exist a Hilbert space ${\mathcal{K}}$, a *-homomorphism $\widehat{\psi}:C^{*}({\mathcal{A}})\to B({\mathcal{K}})$ and a bounded operator $V:{\mathcal{H}}\to{\mathcal{K}}$ such that $\psi(x)=V^{*}\,\widehat{\psi}(x)\,V$ for all $x\in C^{*}({\mathcal{A}})$. Let $\widehat{\pi}=(\widehat{\pi}_{g})_{g\in G}$ denote the representation of ${\mathcal{A}}$ on ${\mathcal{K}}$ associated to $\widehat{\psi}$, so that
@@
By Fell’s absorption principle for Fell bundles [28, Proposition 18.4], the integrated forms $\phi_{\mathcal{A}}$ and $\phi_{\mathcal{B}}$ of the representations $(\widehat{\pi}_{g}\otimes\lambda^{G}_{g})_{g\in G}$ and $(\pi_{h}\otimes\lambda^{H}_{h})_{h\in H}$ (as defined in [28, Prop 16.29]) factor through $C^{*}_{r}({\mathcal{A}})$ and $C^{*}_{r}({\mathcal{B}})$, respectively. This means that there are *-homomorphisms $\psi_{\mathcal{A}}:C^{*}_{r}({\mathcal{A}})\to B({\mathcal{K}}\otimes\ell^{2}(G))$ and $\psi_{\mathcal{B}}:C^{*}_{r}({\mathcal{B}})\to B({\mathcal{H}}\otimes\ell^{2}(H))$ such that $\phi_{\mathcal{A}}=\psi_{\mathcal{A}}\circ\Lambda^{\mathcal{A}}$ and $\phi_{\mathcal{B}}=\psi_{\mathcal{B}}\circ\Lambda^{\mathcal{B}}$. Using [28, Corollary 18.5], we may regard $\psi_{\mathcal{A}}$ as a map from $C^{*}_{r}({\mathcal{A}})$ into $B({\mathcal{K}})\otimes_{\rm min}C_{r}^{*}(G)$ and $\psi_{\mathcal{B}}$ as a map from $C^{*}_{r}({\mathcal{B}})$ into $B({\mathcal{H}})\otimes_{\rm min}C_{r}^{*}(H)$, which satisfy that
@@
for all $g\in G$, $a\in A_{g}$, $h\in H$ and $b\in B_{h}$. Note that $\psi_{\mathcal{B}}$ is faithful since $\pi_{e}$ is assumed to be so, cf. [28, Proposition 18.4].
Now, as shown in [11, Corollary 8.C.15], since $\ker(\varphi)$ is amenable, $\varphi$ extends to a *-homomorphism $\varphi_{*,r}:C_{r}^{*}(G)\to C_{r}^{*}(H)$ satisfying that $\varphi_{*,r}(\lambda_{g}^{G})=\lambda^{H}_{\varphi(g)}$ for all $g\in G$.
@@
$~\Phi(\lambda^{\mathcal{A}}_{g}(a))$ $=(\widetilde{V}\otimes_{\rm min}\varphi_{*,r})(\psi_{\mathcal{A}}(\lambda^{\mathcal{A}}_{g}(a)))=(\widetilde{V}\otimes_{\rm min}\varphi_{*,r})(\widehat{\pi}_{g}(a)\otimes\lambda^{G}_{g})$
$=V^{*}\widehat{\pi}_{g}(a)V\otimes\lambda^{H}_{\varphi(g)}=V^{*}\widehat{\psi}(\widehat{j}^{\mathcal{A}}_{g}(a))V\otimes\lambda^{H}_{\varphi(g)}$
@@
Since $\psi_{\mathcal{B}}$ is faithful, we can define $\psi^{\prime}:C_{r}^{*}({\mathcal{A}})\to C_{r}^{*}({\mathcal{B}})$ by $\psi^{\prime}=(\psi_{\mathcal{B}})^{-1}\circ\Phi$, which then satisfies $\psi^{\prime}(\lambda^{\mathcal{A}}_{g}(A_{g}))\subseteq\lambda^{\mathcal{B}}_{\varphi(g)}(B_{\varphi(g)})$, i.e., $\psi^{\prime}$ is $\varphi$-diagonal. It is now straightforward to check that $\psi^{\prime}$ also satisfies (3). $\square$
@@
for $g\in G$ and $a\in A_{g}$. Since $\pi_{\varphi(g)}$ is injective on $B_{\varphi(g)}$ (this readily follows from the definition of a *-representation [28, Def. 16.20] and the fact that $\pi_{e}$ is faithful by assumption), this implies that
$\phi^{\prime}(\lambda^{\mathcal{A}}_{g}(a))=\lambda^{\mathcal{B}}_{\varphi(g)}(T_{g}(a))\quad\text{for every }g\in G\text{ and }a\in A_{g}.$
@@
Even if ${\rm ker}(\varphi)$ is not amenable it may still happen that a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map $T$ is reduced with $M_{T}$ completely positive, cf. Example 4.1. Hence, in such a setting, it would be interesting to find some additional condition(s) ensuring that $T$ is reduced with $M_{T}$ completely positive.
@@
Given a completely positive linear map $\Phi:C^{*}({\mathcal{A}})\to C^{*}({\mathcal{B}})$, let $T^{\Phi}$ denote the positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map associated to it in Proposition 3.9. Then, by Theorem 3.11, the map $\Phi_{T^{\Phi}}$ is a $\varphi$-diagonal completely positive linear map (which could be called the $\varphi$-diagonalization of $\Phi$). Hence the map $\Phi\mapsto\Phi_{T^{\Phi}}$ is a projection from the space CP($C^{*}({\mathcal{A}}),C^{*}({\mathcal{B}})$) of completely positive linear maps onto the subspace of $\varphi$-diagonal maps. If the kernel of $\varphi$ is amenable, using Proposition 3.8 and Theorem 3.14, one obtains a similar projection map on $CP(C_{r}^{*}({\mathcal{A}}),C_{r}^{*}({\mathcal{B}}))$.
@@
Then any $S\in B({\mathcal{A}},\varphi,{\mathcal{B}})$ is a full ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map, which is also reduced if ${\rm ker}(\varphi)$ is assumed to be amenable. This follows readily from Theorems 3.11 and 3.14 after noticing that the map $T\mapsto\phi_{T}$ is linear. We thus get a linear map $T\mapsto\Phi_{T}$ from $B({\mathcal{A}},\varphi,{\mathcal{B}})$ into the space of completely bounded linear maps CB($C^{*}({\mathcal{A}}),C^{*}({\mathcal{B}})$), and a similar one into CB($C_{r}^{*}({\mathcal{A}}),C_{r}^{*}({\mathcal{B}})$) when ${\rm ker}(\varphi)$ is amenable. It is easy to see that the space $B({\mathcal{A}}):=B({\mathcal{A}},{\rm id}_{G},{\mathcal{A}})$ becomes a unital algebra w.r.t. natural composition, which could be called the Fourier-Stieltjes algebra of ${\mathcal{A}}$, in analogy with the case where ${\mathcal{A}}$ is the Fell bundle associated to a unital discrete twisted $C^{*}$-dynamical system, cf. [10].
@@
Remark 3.19.
One can define a category ${\tt Fell}$-${\tt PD}$ whose objects are Fell bundles over discrete groups, and arrows from ${\mathcal{A}}=(A_{g})_{g\in G}$ to ${\mathcal{B}}=(B_{h})_{h\in H}$ are pairs $(\varphi,T)$, where $\varphi:G\to H$ is a homomorphism and $T$ is a positive definite ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map, composition of arrows being defined in the natural way.
We then get a functor $\mathcal{F}$ from the category ${\tt Fell}$-${\tt PD}$ to the category ${\tt C^{*}CP}$ whose objects are $C^{*}$-algebras and arrows are completely positive linear maps, by setting